$SPAD/src/input d03edf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)cl all
showArrayValues true
showScalarValues true
ngx := 9
ngy := 9
lda :=  134
alpha := 1.7
hx := 1/(ngx+1)
hy := 1/(ngy+1)
a := new(lda,7,0.0)$Matrix DoubleFloat;
rhs := new(1,lda,0.0)$Matrix DoubleFloat;
ub := new(1,ngx*ngy,0.0)$Matrix DoubleFloat;
for j in 1..ngy repeat
  for i in 1..ngx repeat
    k := (j-1)*ngx + i
    a(k,1) := 1 - 0.5*alpha
    a(k,2) := 0.5*alpha
    a(k,3) := 1 - 0.5*alpha
    a(k,4) := -4 + alpha
    a(k,5) := 1 - 0.5*alpha
    a(k,6) := 0.5*alpha
    a(k,7) := 1 - 0.5*alpha
    rhs(1,k) := -4.0*hx*hy
for i in 2..(ngx-1) repeat
  ix := i
  a(ix,1) := 0
  a(ix,2) := 0
  ix := i + (ngy -1)*ngx
  a(ix,6) := 0
  a(ix,7) := 0
for j in 2..(ngy-1) repeat
  iy := (j-1)*ngx+1
  a(iy,3) := 0
  a(iy,6) := 0
  iy := j *ngx
  rhs(1,iy) := rhs(1,iy) - a(iy,5) - a(iy,2)
  a(iy,2) := 0
  a(iy,5) := 0

k := 1
a(1,1) := 0
a(1,2) := 0
a(1,3) := 0
a(1,6) := 0
k := 1 + (ngy-1)*ngx
a(k,3) := 0
a(k,6) := 0
a(k,7) := 0
k := ngx
rhs(1,k) := rhs(1,k) - a(k,2)*0.5 -a(k,5)
a(k,1) := 0
a(k,2) := 0
a(k,5) := 0
k := ngx * ngy
rhs(1,k) := rhs(1,k) - a(k,2) - a(k,5)
a(k,2) := 0
a(k,5) := 0
a(k,6) := 0
a(k,7) := 0
ifail := 0
maxit := 15
acc := 1.0e-4
iout := 0
result:=d03edf(ngx,ngy,lda,maxit,acc,iout,a,rhs,ub,ifail)

———-

References^{[1] nothing}

5